Explain the meaning of the phrase number representation, and give two examples in which the 

number “2” may be represented by your learners in the intermediate phase 




1
Expert's answer
2020-09-15T16:04:42-0400

Phrase number representation is any notation for the representation of numbers. A number representation is a description of what it is made of and what it is. For example, people use numbers to describe how many thing they have.

Let's represent number “2” . Number “2” consists of the sum of units

So we can represent "2" as sum of 1 and 1: "1 + 1 = 2"



We also can represent "2" as difference of 5 and 3: "5 - 3 = 2" ;



division of 6 by 3: "6 \/ 3 = 2"



and multiplication of 0.5 and 4: "0.5 * 4 = 2" :